A Bibliography of Supercomputing ’90

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 February 2022 Version 1.04

Abstract

This bibliography records articles presented at the Supercomputing ’90 conference.

Title word cross-reference

N [BME90].

-body [BME90].

2 [Ess90, MLR90].

3 [Chr90]. 3-D [Chr90]. 3D [TMH^{+}90, WNA90].

600Js [WTW90].

’90 [IEE90].

A-NET [BYI^{+}90]. abstraction [AGNS90]. accessibility [ADT90]. accessibility-based [ADT90]. adequacy [Azm90]. advanced [KHN90]. algebra [ABD^{+}90, Che90, JHJ90]. algorithm [Duf90, Har90, NEK90, RT90, Sto90]. aliasing [HS90]. allocation [AG90, BIY90, HS90]. ALPS [Vis90]. analysis [Che90, FCS90, Har90, HK90, LMR90, MA90, RT90, SF90]. analyzer [HLB90]. annealing [HIJ90]. anomaly [HKMC90]. Application [BW90, HIJ90, HMMP90, MLR90]. applications [CP90, HLB90]. approach [CLS90, Gup90, SN90a, WTW90]. arbitrary [KM90]. Architectural [HS90]. Architecture [CHJ^{+}90, BYI^{+}90, FCS90, Sti90]. architectures [MA90, SHH90, TLS90]. array [HK90, OH90]. aspects [FCS90, HCZ90]. assisted [CLS90]. balancing [XH90]. banyans [SF90]. barrier [BP90]. based [ADT90, GEW90, Map90, WNA90, Zho90, ADT90, CRV90]. BBN [CM90]. behavior [GC90]. Benchmarks [CFN90b]. binding [SN90a]. black [ST90]. Blitz [HN90]. blocking [Che90]. body [BME90]. buffered [SF90]. Building [ST90, SP90].

Cache [WSH90, ADT90, CP90, GC90]. caches [ADT90]. Carlo [CFN90a, HMMP90, IK90, MK90, RT90]. case [Fre90]. Cedar [SMBSS90]. Center [IEE90]. CERFACS [Duf90]. channels [WSH90]. characterization [WMK90]. Chess [HN90, BKN90]. chip [GEW90]. chromodynamics [HMMP90]. chromosomes [WTW90]. cinema [ST90]. Clemson [SP90]. cluster [WTW90]. clustered [KSZ90]. code [CM90, MK90]. coherence [ADT90, WSH90]. communicating [RNB90]. communication [AGNS90, CE90, FCS90, LC90a, WSH90, Zho90]. communication/data [FCS90]. communication/data-conditional [FCS90]. compare [Sto90]. compare-and-swap [Sto90]. comparison [CF90a, SW90]. compiler [CLS90]. compiler-assisted [CLS90]. composites [LCD90]. Computation [NEK90]. Computational [SP90, PC90, Sti90, WTW90]. computations [FCS90]. Computer [LCD90, CM90, Zho90]. computers [ABD^{+}90, BIY90, CBZ90, CF90b, WG90]. computing [AGNS90, KSZ90]. concepts [BKN90]. concurrent [SMBSS90]. conditional [FCS90]. conformational [RT90]. connected [CE90]. Connection [BCK90, BME90, CC90]. constrained [NEK90]. control [KM90]. coordination [LS90]. coprocessor [HB90]. correct [Sto90]. Cost [MA90, Nat90]. Cost-performance [MA90]. course [Han90]. CRAY [WMK90, Ess90, HN90, MLR90]. Cray-2 [Ess90]. crossbars [SF90]. cut [SF90]. cut-through [SF90].

D [Chr90]. Data

[CP90, Che90, HCZ90, HMMP90, LMR90]. data-conditional [FCS90]. Davis [CF90b]. Davis-Putman [CF90b]. debugging [HKMC90]. declarative [RNB90]. decomposition [CFN90a, CFN90b, Har90]. Delirium [LS90]. dependence [LC90b]. Design [MTSA90, SF90, GEW90]. details [HN90]. detection [HKMC90]. device [TMH^{+}90]. dilation [HTCD90]. direct [BME90]. directional [CD90]. displacement [Gup90]. distributed [AG90, HIJ90, HB90]. distribution [KM90]. dynamic [XH90].

effects [HK90]. Efficient [CC90, CFN90a]. electron [MK90]. element [Chr90]. Embedding [RWY90, LS90]. environments [Map90]. equilibration [NEK90]. ES/3090 [WTW90]. estimation [AC90]. evaluating [BKN90]. evaluation [CE90, SW90]. execution [CLS90, Gup90, MLR90]. exits [TLS90]. Experience [HLB90, HK90]. Experiences [SP90]. Experimental [FCS90]. explicit [LC90a].

factorization [RG90]. Fast [BP90, Nat90]. Fault [CRV90]. Fault-tolerant [CRV90]. FFT [CFN90a]. fiber [Szy90]. Fine [Vis90, KHN90]. Fine-grain [Vis90]. finite [Chr90]. flexible [AGNS90]. flow [Che90, KM90, LMR90]. flows [ST90]. fluid [ST90]. fly [HKMC90]. Folding [Ess90]. formal [HCZ90]. format [WNA90]. Fortran [CF90a]. Future [SHH90].

general [SHH90, WNA90]. generalized [CF90b]. Generating [LC90a]. generation [IK90, WG90]. generator [FSV90]. geometric [HCZ90]. grain [KHN90, Vis90]. graph [RWY90]. graphs [Sti90]. gravitational [ST90]. Group [Sti90].

hardware [BP90]. heat [Chr90, MK90]. heterogeneity [MA90]. Heuristic [XH90]. Hierarchical [Che90]. High [TMH^{+}90, ABD^{+}90, Cli90, HMMP90, LCD90, Map90, WG90]. high-performance [ABD^{+}90, Map90]. high-speed [LCD90, WG90]. high-temperature [HMMP90]. Hilton [IEE90]. holes [ST90]. hypercube [AG90, BME90]. hypercubes [CD90]. hypermesh [Szy90].

IBM [SW90, WTW90]. impact [HH90, LCD90]. Implementation [RSV90, CHJ^{+}90]. Improving [GC90, RG90]. independent [BIY90]. Information [HMMP90]. instruction [GC90]. interconnection [Map90]. interface [Chr90]. interleaved [RH90]. interprocedural [HK90]. interprocessor [CE90]. introduction [Han90, HCZ90]. Ising [IK90].

language [BCK90, LS90, Vis90]. LAPACK [ABD^{+}90]. Large [KSZ90, AG90, NEK90]. level [MTSA90]. library [ABD^{+}90]. light [PC90]. linear [ABD^{+}90, Che90, OH90]. Lisp [WNA90]. Lisp-based [WNA90]. Livermore [CF90a]. load [XH90]. Logarithmic [Nat90]. logic [CF90b, CC90]. long [FSV90]. long-period [FSV90]. Loop [Gup90, KM90]. loops [CH90, Gup90, TLS90, CF90a].

machine [RSV90, BCK90, BME90, CC90]. machines [Szy90]. many [WSH90]. mapping [WTW90]. massive [LC90b]. Massively [PC90, AC90, Sti90]. matrix [HH90, NEK90, RG90]. memories [RH90]. memory [AGNS90, HH90, HB90, LC90a, Map90, OH90]. memory-based [Map90]. mesh [CE90]. mesh-connected [CE90]. meshes [RWY90]. message [Azm90, HB90, XH90]. message-passing [Azm90, XH90]. methods [PC90, XH90]. microtasking [CH90]. MIMD [Szy90]. MIN [ADT90, CRV90]. MIN-based [ADT90, CRV90]. MINIMOS [TMH^{+}90]. mixed [FCS90]. mixed-mode [FCS90]. mode [FCS90]. model [Chr90, IK90, Zho90]. modeling [BW90, HCZ90, HIJ90]. models [WNA90]. molecular [HIJ90]. monitor [WMK90]. monitoring [SMBSS90]. MONT3E [MK90]. Monte [MK90, CFN90a, HMMP90, IK90, RT90]. moving [Chr90]. MP [MLR90, WMK90]. multicomputer [Map90]. multicomputers [CE90, HB90]. Multilinear [JHJ90]. multiplication [HH90]. multiprocessor [GEW90, KHN90, RG90, SMBSS90]. multiprocessors [ADT90, CFN90b, KSZ90]. multithreaded [HLB90].

near [KHN90]. nested [CH90]. NET [BYI^{+}90]. network [BIY90, Fre90]. network-topology [BIY90]. networking [Cli90]. networks [CE90]. neutron [Azm90]. nonshared [AGNS90]. November [IEE90]. number [FSV90, IK90]. numerical [Che90].

object [BYI^{+}90]. object-oriented [BYI^{+}90]. on-the-fly [HKMC90]. optic [Szy90]. optimal [BME90, Nat90]. optimally [KHN90]. optimization [HMMP90]. optimized [SW90]. optimizing [HN90]. organization [HH90]. oriented [BYI^{+}90]. OSCAR [KHN90].

P3D [WNA90]. paradigm [RSV90]. Parallel [Duf90, HKMC90, KHN90, AC90, AGNS90, Azm90, BIY90, BYI^{+}90, CM90, CH90, CC90, CFN90a, FCS90, Gup90, JHJ90, LMR90, MTSA90, Nat90, OH90, PC90, RSV90, RT90, SN90a, Sti90, SN90b, WTW90, WSH90, Zho90]. parallelism [LC90b, Vis90]. Parallelization [CM90, TLS90]. Parameterized [Har90]. partially [OH90]. particles [PC90]. Partitioning [RNB90]. passing [Azm90, HB90, XH90]. PDE [CFN90a, OH90]. Perfect [CFN90b]. Performance [AC90, CE90, SW90, ABD^{+}90, CP90, Cli90, CFN90a, CFN90b, HH90, HLB90, Har90, Map90, MA90, RG90, TMH^{+}90, WMK90]. Performing [LMR90]. period [FSV90]. petroleum [BW90]. phase [Chr90]. physical [WTW90]. pipelined [TLS90]. pollution [GC90]. portable [ABD^{+}90]. practice [Nat90]. preconditioning [TMH^{+}90]. presence [HS90]. primitives [CBZ90]. problems [Azm90, NEK90]. procedure [CF90b, Zho90]. procedure-based [Zho90]. Proceedings [IEE90]. processes [RNB90]. processing [HIJ90, KHN90]. processor [IK90, SW90]. processors [OH90, SW90, WSH90]. program [HKMC90, LC90a, MLR90, SP90]. programming [JHJ90, RSV90, SN90a, Vis90]. programs [RNB90, SMBSS90]. propositional [CF90b]. proteins [RT90]. protocol [ADT90]. proving [CF90b]. purpose [SHH90]. Putman [CF90b].

Quantum [HIJ90, HMMP90]. queue [Sto90].

radiation [CM90]. random [FSV90, IK90]. randomly [RH90]. real [Han90]. reducing [GC90]. references [LC90a]. register [FSV90, HS90]. relativistic [ST90]. relativity [HTCD90]. rendering [HCZ90]. representing [WNA90]. research [Duf90]. reservoir [BW90]. Resource [SN90a]. response [LCD90]. RISC [GEW90, SW90]. RNA [Ess90]. Rockefeller [IEE90]. routed [CE90]. routing [CRV90]. Run [SMBSS90]. Run-time [SMBSS90].

scalable [MTSA90]. scalar [SW90]. scale [KSZ90, NEK90]. Scan [CBZ90]. scattering [PC90]. scheduled [KHN90]. scheduling [Gup90, KHN90]. scheme [CH90]. Sciences [SP90]. Scientific [HCZ90, WMK90]. search [RT90]. selection [Fre90]. semi [AG90]. sequential [BW90]. shared [LC90a, OH90, Sto90]. shared-memory [LC90a]. shared-queue [Sto90]. shift [FSV90]. shift-register [FSV90]. side [HK90]. SIMD [Szy90]. SIMD/MIMD [Szy90]. simple [Sto90]. simplex [CFN90a]. simulated [HIJ90]. simulation [CM90, Chr90, CC90, IK90, LCD90]. simulations [CFN90a]. simulator [MTSA90, TMH^{+}90]. SISAL [CF90a]. small [PC90]. solution [OH90]. solver [BME90]. solvers [CFN90a]. solving [Azm90]. sorting [Nat90]. Soviet [WG90]. sparse [CFN90a, RG90]. speed [LCD90, WG90]. speedup [SN90b]. split [ADT90]. splitting [NEK90]. SPMD [CLS90]. stacks [CH90]. staging [BW90]. star [RWY90]. static [KHN90]. strategy [AG90, BIY90]. strength [BKN90]. study [Fre90]. Subdomain [LC90b]. Supercomputer [Fre90, CP90, CHJ^{+}90, MA90, ST90, SHH90, XH90]. supercomputers [AG90, Azm90, CRV90, HN90, TMH^{+}90]. Supercomputing [IEE90, Cli90, Han90]. support [HS90]. supporting [Zho90]. swap [Sto90]. Switch [CH90, MTSA90]. switch-level [MTSA90]. Switch-stacks [CH90]. switching [SF90]. symmetry [Sti90]. synchronization [BP90]. synthetic [FCS90]. system [AC90, Map90, SW90]. system/6000 [SW90]. systems [WSH90].

task [AG90, BIY90]. tasks [BW90, KHN90]. taxonomy [BKN90]. TC2000 [CM90].

Techniques [RG90, CC90]. temperature [HMMP90]. test [LC90b]. Theorem [CF90b]. Time [HTCD90, Nat90, SMBSS90]. tolerant [CRV90]. topology [BIY90]. total [BYI^{+}90]. Tracing [MLR90]. training [Han90]. transfer [MK90]. transformation [Chr90]. transforming [Gup90]. transient [Chr90]. transport [Azm90, CM90, Chr90]. two [Chr90, SW90, WMK90]. two-phase [Chr90].

UC [BCK90]. Uni [CD90]. Uni-directional [CD90]. universal [SN90a]. update [ADT90]. user [Han90]. using [CF90a, CF90b, KHN90, Sto90, WTW90, WMK90].

VAX [CFN90b]. vector [CBZ90, CF90b, IK90, KSZ90, SW90]. vectorized [Chr90, FSV90]. versus [CF90a]. Very [Cli90]. via [FCS90]. view [SN90b]. visualization [HCZ90, HIJ90, HTCD90]. VLIW [CHJ^{+}90]. VLSI [MTSA90].

waves [ST90]. workloads [WMK90]. workstations [RG90]. wormhole [CE90]. wormhole-routed [CE90]. write [ADT90].

X [MLR90, WMK90]. X-MP [MLR90, WMK90].

York [IEE90].

References

Angerson:1990:LPL

[ABD^{+}90] E. Angerson, Z. Bai, J. Dongarra, A. Greenbaum, A. McKenney, J. Du Croz, S. Hammarling, J. Demmel, C. Bischof, and D. Sorensen. LAPACK: A portable linear algebra library for high-performance computers. In IEEE [IEE90], pages 2–11. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Agrawal:1990:PEM

[AC90] V. D. Agrawal and S. T. Chakradhar. Performance estimation in a massively parallel system. In IEEE [IEE90], pages 306– 313. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Algudady:1990:WUC

[ADT90] M. S. Algudady, C. R. Das, and M. J. Thazhuthaveetil. A write update cache coherence protocol for MIN-based multiprocessors with accessibility-based split caches. In IEEE [IEE90], pages 544–553. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Ahmad:1990:SDT

[AG90] I. Ahmad and A. Ghafoor. A semi distributed task allocation strategy for large hypercube supercomputers. In IEEE [IEE90], pages 898–907. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Alverson:1990:FCA

[AGNS90] G. A. Alverson, W. G. Griswold, D. Notkin, and L. Snyder. A flexible communication abstraction for nonshared memory parallel computing. In IEEE [IEE90], pages 584–593. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Azmy:1990:AMP

[Azm90] Y. Y. Azmy. On the adequacy of message-passing parallel supercomputers for solving neutron transport problems. In IEEE [IEE90], pages 693–699. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Bagrodia:1990:ULC

[BCK90] R. Bagrodia, K. M. Chandy, and E. Kwan. UC: a language for the Connection Machine. In IEEE [IEE90], pages 525–534. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Baba:1990:NIT

[BIY90] T. Baba, Y. Iwamoto, and T. Yoshinaga. A networktopology independent task allocation strategy for parallel computers. In IEEE [IEE90], pages 878–887. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Berliner:1990:TCE

[BKN90] H. Berliner, D. Kopec, and E. Northam. A taxonomy of concepts for evaluating chess strength. In IEEE [IEE90], pages 336–343. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Brunet:1990:OHD

[BME90] J.-P. Brunet, J. P. Mesirov, and A. Edelman. An optimal hypercube direct N-body solver on the Connection Machine. In IEEE [IEE90], pages 748–752. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Beckmann:1990:FBS

[BP90] C. J. Beckmann and C. D. Polychronopoulos. Fast barrier synchronization hardware. In IEEE [IEE90], pages 180–189. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Briens:1990:ASS

[BW90] F. J. L. Briens and C. H. Wu. Application of sequential staging of tasks to petroleum reservoir modeling. In IEEE [IEE90], pages 428–435. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Baba:1990:POT

[BYI^{+}90] T. Baba, T. Yoshinaga, T. Iijima, Y. Iwamoto, Hamada, M., and M. Suzuki. A parallel objectoriented total architecture: A-NET. In IEEE [IEE90], pages 276–285. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chatterjee:1990:SPV

[CBZ90] S. Chatterjee, G. E. Blelloch, and M. Zagha. Scan primitives for vector computers. In IEEE [IEE90], pages 666–675. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chung:1990:EPL

[CC90] M. J. Chung and Y. Chung. Efficient parallel logic simulation techniques for the Connection Machine. In IEEE [IEE90], pages 606–614. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chou:1990:UH

[CD90] C.-H. Chou and D. H. C. Du. Unidirectional hypercubes. In IEEE [IEE90], pages 254–263. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chittor:1990:PEM

[CE90] S. Chittor and R. Enbody. Performance evaluation of meshconnected wormhole-routed networks for interprocessor communication in multicomputers. In IEEE [IEE90], pages 647– 656. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Cann:1990:SVF

[CF90a] D. Cann and J. Feo. SISAL versus Fortran: a comparison using the Livermore Loops. In IEEE [IEE90], pages 626–636. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chen:1990:TPP

[CF90b] Wen-Tsuen Chen and Ming-Yi Fang. Theorem proving in propositional logic on vector computers using a generalized Davis-Putman procedure. In IEEE [IEE90], pages 658–665. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Cvetanovic:1990:EDP

[CFN90a] Z. Cvetanovic, E. G. Freedman, and C. Nofsinger. Efficient decomposition and performance of parallel PDE, FFT, Monte Carlo simulations, simplex, and sparse solvers. In IEEE [IEE90], pages 465–474. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Cvetanovic:1990:PBD

[CFN90b] Z. Cvetanovic, E. G. Freedman, and C. Nofsinger. Perfect Benchmarks decomposition and performance on VAX multiprocessors. In IEEE [IEE90], pages 455–464. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chow:1990:SSM

[CH90] J.-H. Chow and W. L. Harrison, III. Switch-stacks: A scheme for microtasking nested parallel loops. In IEEE [IEE90], pages 190–199. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback:

ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chen:1990:HBD

[Che90] D. Chen. Hierarchical blocking and data flow analysis for numerical linear algebra. In IEEE [IEE90], pages 12–19. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Colwell:1990:AIV

[CHJ^{+}90] R. P. Colwell, W. E. Hall, C. S.

Joshi, D. B. Papworth, P. K. Rodman, and J. E. Tornes. Architecture and implementation of a VLIW supercomputer. In IEEE [IEE90], pages 910–919. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Christon:1990:VFE

[Chr90] M. Christon. A vectorized 3-D finite element model for transient simulation of two-phase heat transport with phase transformation and a moving interface. In IEEE [IEE90], pages 436– 445. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Clinger:1990:VHP

[Cli90] M. Clinger. Very high performance networking for supercomputing. In IEEE [IEE90], pages 160–168. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Cytron:1990:CAS

[CLS90] R. Cytron, J. Lipkis, and E. Schonberg. A compiler-assisted approach to SPMD execution. In IEEE [IEE90], pages 398– 406. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Celmaster:1990:PRT

[CM90] W. Celmaster and E. N. May. Parallelization of a radiation transport simulation code on the BBN TC2000 parallel computer. In IEEE [IEE90], pages 448– 454. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback:

ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Callahan:1990:DCP

[CP90] D. Callahan and A. Porterfield.

Data cache performance of supercomputer applications. In IEEE [IEE90], pages 564–572. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chalasani:1990:FRM

[CRV90] S. Chalasani, C. S. Raghavendra, and A. Varma. Fault-tolerant routing in MIN-based supercomputers. In IEEE [IEE90], pages 244–253. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Duff:1990:PAR

[Duf90] I. S. Duff. Parallel algorithm research at CERFACS. In IEEE [IEE90], pages 536–542. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056.

IEEE catalog number 90CH2916-5.

Ess:1990:FRC

[Ess90] M. Ess. Folding RNA on the Cray-2. In IEEE [IEE90], pages 103–111. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Fineberg:1990:EAC

[FCS90] S. A. Fineberg, T. L. Casavant, and H. J. Siegel. Experimental analysis of communication/dataconditional aspects of a mixedmode parallel architecture via synthetic computations. In IEEE [IEE90], pages 637–646. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

French:1990:SNS

[Fre90] W. L. French. Supercomputer network selection: A case study. In IEEE [IEE90], pages 154– 159. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Filippone:1990:VLS

[FSV90] Salvatore Filippone, Paolo Santangelo, and Marcello Vitaletti. A vectorized long-period shiftregister random number generator. In IEEE [IEE90], pages 676– 684. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Gupta:1990:IIC

[GC90] R. Gupta and C.-H. Chi. Improving instruction cache behavior by reducing cache pollution. In IEEE [IEE90], pages 82–91. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Gupta:1990:DRB

[GEW90] R. Gupta, M. Epstein, and M. Whelan. The design of a RISC based multiprocessor chip. In IEEE [IEE90], pages 920– 929. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Gupta:1990:LDA

[Gup90] R. Gupta. Loop displacement: an approach for transforming and scheduling loops for parallel execution. In IEEE [IEE90], pages 388–397. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hanson:1990:RIS

[Han90] F. B. Hanson. A real introduction to supercomputing: a user training course. In IEEE [IEE90], pages 376–385. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Harkin:1990:PAD

[Har90] G. J. Harkin. Parameterized algorithm decomposition and performance analysis. In IEEE [IEE90], pages 314–323. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hsu:1990:MPC

[HB90] J.-M. Hsu and P. Banerjee. A message passing coprocessor for distributed memory multicomputers. In IEEE [IEE90], pages 720–729. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Harrand:1990:SDV

[HCZ90] V. J. Harrand, A. Choudry, and J. P. Ziebarth. Scientific data visualization: a formal introduction to the rendering and geometric modeling aspects. In IEEE [IEE90], pages 775–783. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hake:1990:IMO

[HH90] J.-F. Hake and W. Homberg. The impact of memory organization on the performance of matrix multiplication. In IEEE [IEE90], pages 34–40. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hohl:1990:QMM

[HIJ90] D. Hohl, R. Idaszak, and R. O.

Jones. Quantum molecular modeling with simulated annealing — a distributed processing and visualization application. In IEEE [IEE90], pages 816–825. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Havlak:1990:EIA

[HK90] P. Havlak and K. Kennedy. Experience with interprocedural analysis of array side effects. In IEEE [IEE90], pages 952–961. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hood:1990:PPD

[HKMC90] R. Hood, K. Kennedy, and J. Mellor-Crummey. Parallel program debugging with on-the-fly anomaly detection. In IEEE [IEE90], pages 74–81. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hansen:1990:EPA

[HLB90] G. J. Hansen, C. A. Linthicum, and G. Brooks. Experience with a performance analyzer for multithreaded applications. In IEEE [IEE90], pages 124–131. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Huang:1990:IOM

[HMMP90] S. Huang, K. J. M. Moriarty, E. A. Myers, and J. Potvin. Information optimization for Monte Carlo data and application to high-temperature quantum chromodynamics. In IEEE [IEE90], pages 742–747. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hyatt:1990:CSD

[HN90] R. M. Hyatt and H. L. Nelson.

Chess and supercomputers: details about optimizing Cray Blitz. In IEEE [IEE90], pages 354– 363. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Heggy:1990:ASR

[HS90] B. Heggy and M. L. Soffa. Architectural support for register allocation in the presence of aliasing. In IEEE [IEE90], pages 730– 739. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Hsiung:1990:TDV

[HTCD90] P.-K. Hsiung, R. H. Thibadeau, C. B. Cox, and R. H. P. Dunn. Time dilation visualization in relativity. In IEEE [IEE90], pages 835–844. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

IEEE:1990:PSN

[IEE90] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Ito:1990:MCS

[IK90] N. Ito and Y. Kanada. Monte Carlo simulation of the Ising model and random number generation on the vector processor. In IEEE [IEE90], pages 753– 763. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Johnson:1990:MAP

[JHJ90] R. W. Johnson, C.-H. Huang, and J. R. Johnson. Multilinear algebra and parallel programming. In IEEE [IEE90], pages 20–31. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Kasahara:1990:PPN

[KHN90] H. Kasahara, H. Honda, and S. Narita. Parallel processing of near fine grain tasks using static scheduling on OSCAR (optimally scheduled advanced multiprocessor). In IEEE [IEE90], pages 856–864. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903.

IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Kennedy:1990:LDA

[KM90] K. Kennedy and K. S. McKinley. Loop distribution with arbitrary control flow. In IEEE [IEE90], pages 407–416. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Kamel:1990:LSC

[KSZ90] A. Kamel, P. Sguazzero, and V. Zecca. Large scale computing on clustered vector multiprocessors. In IEEE [IEE90], pages 418–427. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Li:1990:GEC

[LC90a] J. Li and M. Chen. Generating explicit communication from shared-memory program references. In IEEE [IEE90], pages 865–876. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Lu:1990:SDT

[LC90b] L.-C. Lu and M. C. Chen. Subdomain dependence test for massive parallelism. In IEEE [IEE90], pages 962–972. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Langlie:1990:CSH

[LCD90] S. Langlie, W. Cheng, and I. Dilber. Computer simulation of highspeed impact response of composites. In IEEE [IEE90], pages 826–834. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Lee:1990:PDF

[LMR90] Y.-F. Lee, T. J. Marlowe, and B. G. Ryder. Performing data flow analysis in parallel. In IEEE [IEE90], pages 942–951. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-

Lucco:1990:DEC

[LS90] S. Lucco and O. Sharp. Delirium: an embedding coordination language. In IEEE [IEE90], pages 515–524. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Menasce:1990:CAH

[MA90] D. Menasce and V. Almeida. Cost-performance analysis of heterogeneity in supercomputer architectures. In IEEE [IEE90], pages 169–177. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Maples:1990:HMI

[Map90] C. Maples. A high-performance, memory-based interconnection system for multicomputer environments. In IEEE [IEE90], pages 295–304. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Maltby:1990:MMC

[MK90] J. D. Maltby and B. T. Kornblum.

MONT3E: A Monte Carlo electron heat transfer code. In IEEE [IEE90], pages 700–707. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Malony:1990:TAP

[MLR90] A. D. Malony, J. L. Larson, and D. A. Reed. Tracing application program execution on the Cray X-MP and Cray 2. In IEEE [IEE90], pages 60–73. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Mueller-Thuns:1990:DSP

[MTSA90] R. B. Mueller-Thuns, D. G. Saab, and J. A. Abraham. Design of a scalable parallel switch-level simulator for VLSI. In IEEE [IEE90], pages 615–624. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Natvig:1990:LTC

[Nat90] L. Natvig. Logarithmic time cost optimal parallel sorting is not yet fast in practice. In IEEE [IEE90], pages 486–494. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Nagurney:1990:CLS

[NEK90] A. Nagurney, A. Eydeland, and D.-S. Kim. Computation of large scale constrained matrix problems: the splitting equilibration algorithm. In IEEE [IEE90], pages 214–223. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Okawa:1990:LAP

[OH90] Y. Okawa and N. Haraguchi. A linear array of processors with partially shared memory for parallel solution of PDE. In IEEE [IEE90], pages 41–48. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Potter:1990:MPC

[PC90] D. J. Potter and M. P. Cline. Massively parallel computational methods in light scattering by small particles. In IEEE [IEE90], pages 686–692. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Rothberg:1990:TIP

[RG90] E. Rothberg and A. Gupta. Techniques for improving the performance of sparse matrix factorization on multiprocessor workstations. In IEEE [IEE90], pages 232–241. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Raghavan:1990:RIM

[RH90] R. Raghavan and J. P. Hayes.

On randomly interleaved memories. In IEEE [IEE90], pages 49–58. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Roy:1990:PDP

[RNB90] J. M. A. Roy, M. D. Nagel, and L. Bic. Partitioning declarative programs into communicating processes. In IEEE [IEE90], pages 846–855. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Rao:1990:IMP

[RSV90] M. Rao, Z. Segall, and D. Vrsalovic. Implementation machine paradigm for parallel programming. In IEEE [IEE90], pages 594–603. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Ripoll:1990:PMC

[RT90] D. R. Ripoll and S. J. Thomas. A parallel Monte Carlo search algorithm for the conformational analysis of proteins. In IEEE [IEE90], pages 94–102. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Ranka:1990:EMS

[RWY90] S. Ranka, J. Wang, and N. Yeh. Embedding meshes on the star graph. In IEEE [IEE90], pages 476–485. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Szymanski:1990:DAB

[SF90] T. Szymanski and C. Fang. Design and analysis of buffered crossbars and banyans with cutthrough switching. In IEEE [IEE90], pages 264–273. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Smith:1990:FGP

[SHH90] J. E. Smith, W.-C. Hsu, and C. Hsiung. Future general purpose supercomputer architectures. In IEEE [IEE90], pages 796–804. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Sharma:1990:RMC

[SMBSS90] S. Sharma, A. D. Malony, M. W.

Berry, and P. Sinvhal-Sharma. Run-time monitoring of concurrent programs on the Cedar multiprocessor. In IEEE [IEE90], pages 784–793. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Shing:1990:RBU

[SN90a] H. Shing and L. M. Ni. Resource binding — a universal approach to parallel programming. In IEEE [IEE90], pages 574–583. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Sun:1990:AVP

[SN90b] X.-H. Sun and L. M. Ni. Another view on parallel speedup. In IEEE [IEE90], pages 324–333. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Stevenson:1990:EBC

[SP90] D. E. Stevenson and R. M. Panoff. Experiences in building the Clemson Computational Sciences Program. In IEEE [IEE90], pages 366–375. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Shapiro:1990:BBH

[ST90] S. L. Shapiro and S. A. Teukolsky. Building black holes, gravitational waves, and relativistic fluid flows: supercomputer cinema. In IEEE [IEE90], pages 805–814. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Stiller:1990:GGC

[Sti90] L. Stiller. Group graphs and computational symmetry on massively parallel architecture. In IEEE [IEE90], pages 344–353. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Stone:1990:SCS

[Sto90] J. M. Stone. A simple and correct shared-queue algorithm using compare-and-swap. In IEEE [IEE90], pages 495–504. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056.

IEEE catalog number 90CH2916-5.

Simmons:1990:PEI

[SW90] M. L. Simmons and H. J. Wasserman. Performance evaluation of the IBM RISC system/6000: comparison of an optimized scalar processor with two vector processors. In IEEE [IEE90], pages 132–141. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Szymanski:1990:FOH

[Szy90] T. Szymanski. A fiber optic hypermesh for SIMD/MIMD machines. In IEEE [IEE90], pages 710–719. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Tirumalai:1990:PLE

[TLS90] P. Tirumalai, M. Lee, and M. Schlansker. Parallelization of loops with exits on pipelined architectures. In IEEE [IEE90], pages 200–212. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Traar:1990:HPP

[TMH^{+}90] K. P. Traar, W. Mader, O. Heinreichsberger, Selberherr, S., and M. Stiftinger. High performance preconditioning on supercomputers for the 3D device simulator MINIMOS. In IEEE [IEE90], pages 224–231. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Vishnubhotla:1990:FPA

[Vis90] P. Vishnubhotla. Fine-grain parallelism in the ALPS programming language. In IEEE [IEE90], pages 506–514. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Wolcott:1990:SHC

[WG90] P. Wolcott and S. E. Goodman.

Soviet high-speed computers: the new generation. In IEEE [IEE90], pages 930–939. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Williams:1990:CTS

[WMK90] E. Williams, C. T. Myers, and R. Koskela. The characterization of two scientific workloads using the CRAY X-MP performance monitor. In IEEE [IEE90], pages 142–152. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Welling:1990:PLF

[WNA90] J. Welling, C. Nuuja, and P. Andrews. P3D: A Lisp-based format for representing general 3D models. In IEEE [IEE90], pages 766–774. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Willis:1990:CCS

[WSH90] J. C. Willis, A. C. Sanderson, and C. R. Hill. Cache coherence in systems with parallel communication channels and many processors. In IEEE [IEE90], pages 554–563. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903.

IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

White:1990:PCA

[WTW90] S. W. White, D. C. Torney, and C. C. Whittaker. A parallel computational approach using a cluster of IBM ES/3090 600Js for physical mapping of chromosomes. In IEEE [IEE90], pages 112–121. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Xu:1990:HMD

[XH90] J. Xu and K. Hwang. Heuristic methods for dynamic load balancing in a message-passing supercomputer. In IEEE [IEE90], pages 888–897. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Zhou:1990:PCM

[Zho90] J. X. Zhou. A parallel computer model supporting procedurebased communication. In IEEE [IEE90], pages 286–294. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.